# Adaptive resolution of fine scales in modes of microstructured optical fibers

### v1

@@ 5. Three microstructured fibers
5.1. A hollow-core anti-resonant fiber
First, we consider the ARF microstructure in Figure 1(b), motivated by studies in [15, 23] of anti-resonant nodeless tube-lattice fiber. This is fiber has a hollow-core surrounded by capillary glass membranes stuck to a glass cladding.
(a) Initial mesh for ARF geometry.
(b) Final mesh for the ARF, after twenty-nine refinements.
@@ 5.1. A hollow-core anti-resonant fiber
Figure 5. Results from the adaptive algorithm applied to ARF.
@@
[…]
Figure 6. History of adaptive convergence for the ARF fiber.
[…]
Figure 7. History of the confinement loss for the ARF fiber, during the adaptive mesh refinement.
The apparent convergence history during the adaptive process is displayed in Figure 6. Since we do not have exact eigenvalues for this case, we cannot calculate the eigenvalue errors in each iteration. Instead, we compare the Hausdorff distance between the final output $\varLambda_{h}^{\text{final}}$ (at the last iteration) with the discrete eigenvalue cluster $\varLambda_{h}^{(\ell)}$ (with two elements converging to the same number) found at the $\ell$th adaptive iteration. The curve exhibits an overall decreasing trend, even if not a monotonic decrease. The CL values at each iteration are compared to the CL value calculated from the final eigenvalue iterate in Figure 7, which illustrates convergence of the CL approximations.
5.2. A nested anti-resonant nodeless fiber
(a) Initial mesh for the NANF geometry.
(b) Final mesh for the NANF fiber, after thirty-four refinements.
@@ 5.2. A nested anti-resonant nodeless fiber
Figure 8. Results from the adaptive algorithm applied to NANF.
@@
We employ the following parameters for the simulation. The length scale is $L=1.5\times 10^{-5}$ m; the operating wavelength is determined by $k=\frac{2\pi}{1.8}\times 10^{6}\leavevmode\nobreak\ \text{m}^{-1}$; the refractive index is $n_{\rm air}=1.00027717$ in the air regions, and $n_{\rm glass}=1.44087350$ in the glass ring. We utilize the following non-dimensionalized radii and thicknesses: the core of the fiber is a circle of radius $r_{\rm core}=1.0$; the outer capillary is determined by an outer radius $r_{\rm cap,\ o}=0.832$, and a thickness of $t_{\rm cap,\ o}=0.028$, and the inner capillary is defined by an inner radius $r_{\rm cap,\ i}=0.4$, and a thickness of $t_{\rm cap,\ i}=0.028$; the cladding has an inner radius $r_{\rm clad}=1.33333333$, an outer radius $r_{\rm buffer}=2.0$, and a thickness $t_{\rm clad}=0.66666667$; the glass ring has an inner radius $r_{\rm inner}=2.692$, an outer radius $r_{\rm outer}=4.35866667$, and a thickness $t_{\rm clad}=1.66666667$; and the PML region starts at $r_{\rm outer}$ and ends on $r_{\rm PML}=7.02533333$. We consider a PML strength (see (14)) of value $\alpha=2.0$.
[…]
Figure 9. History of adaptive convergence for the NANF fiber.
[…]
Figure 10. History of the confinement loss for the NANF fiber, during the adaptive mesh refinement.
The results in Figure 8 confirm this expectation. The initial mesh (Figure 8(a)) is relatively coarse, with small elements used only to conform to the geometry of the thin glass capillaries and their melding into the cladding (as in the ARF case). We make the elements in the core region smaller (with $0.2\lesssim h\lesssim 0.5$), to capture the expected core-localized fundamental mode of this fiber and verify the saturation assumption. Again, the adaptive meshing algorithm found (in Figure 8(b)) that significant refinements were needed in the outer glass cladding; however, it also determined that some refinements were warranted in the hollow-core near larger capillaries that bound the core. The captured fine-scale modal features are visible in Figure 8(d). The convergence of the eigenvalues and CL values during the adaptive process are illustrated in Figures 9 and 10.
5.3. A photonic bandgap fiber
Finally, we consider the PBG microstructured design in Figure 1(d); previously studied in [24, 7, 19]. The lattice arrangement of the dielectric rods in the fiber can be modified to change the guiding properties of the fiber. We base our model parameters on the descriptions in [19].
@@ 5.3. A photonic bandgap fiber
(b) Final mesh for the PBG fiber, after fifteen refinements.
@@
Figure 11. Numerical results for the PBG fiber.
The scale of the PBG fiber is $L=2.88753\times 10^{-6}$ m. The wavelength we consider is $k=\frac{2\pi}{8.25}\times 10^{7}\text{m}^{-1}$. The refractive index of the cladding and the outer region is $n_{\rm cladding}=1.45$, and the refractive index of the tubes is $n_{\rm tube}=1.8$. We have used a lattice with six dielectric rods, each with a (dimensionless) radius of $r_{\rm tube}=0.57142857$, encircling a fiber core of radius $r_{\rm core}=1.0$; The inner radius of the PML region is $r_{\rm outer}=3.80952380$, and the outer radius of the PML region is $r_{\rm PML}=5.80952380$.
[…]
Figure 12. History of adaptive convergence for the PBG fiber.
[…]
Figure 13. History of the confinement loss for the PBG fiber, during the adaptive mesh refinement.
The results are in Figure 11. The initial mesh (Figure 11(a)) is relatively coarse, yet we employed smaller element in the core (with $h\approx 0.25$). For this fiber, as seen from 11(b), the adaptive process targets refinement within the rods that surround the solid dielectric core. We see finer scale ripples within these rods being resolved in Figure 11(d). Again, these are better seen in the longitudinal component, and are barely seen in the total intensity plot (in Figure 11(c)). The convergence of the eigenvalues during the adaptive process depicted in Figure 12 shows a monotonic descent, unlike what was found for the anti-resonant fibers. (The discrete cluster $\varLambda_{h}^{(\ell)}$ here generally had two elements as in the previous cases.) The convergence of CL values is seen in Figure 13.
@@
Acknowledgments
This work was supported in part by AFOSR Grant FA9550-23-1-0103 and NSF Grant 2245077. The work also benefited from activities organized under the auspices of NSF RTG Grant 2136228.

### v2

@@ 5 Three microstructured fibers
5.1 A hollow-core anti-resonant fiber
First, we consider the ARF microstructure in Figure 1(b), motivated by studies in [15, 23] of anti-resonant nodeless tube-lattice fiber. This fiber has a hollow-core surrounded by glass capillaries (modeled as) embedded into a glass cladding.
(a) Initial mesh for the ARF geometry.
(b) Final mesh for the ARF after the adaptive algorithm
@@ 5.1 A hollow-core anti-resonant fiber
$10^{5}$$10^{6}$$10^{-4}$$10^{-3}$$10^{-2}$umber of d.o.f.s at iteration $\ell$$d(\varLambda_{h}^{\text{final}},\varLambda_{h}^{(\ell)})$
(e) History of adaptive convergence.
$10^{5}$$10^{6}$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$umber of d.o.f.s at iteration $\ell$$\text{CL}_{\ell}$
(f) History of computed loss values.
Figure 6: Results from the adaptive algorithm applied to ARF.
@@
The apparent convergence history during the adaptive process is displayed in Figure 6(e). Since we do not have exact eigenvalues for this case, we cannot calculate the eigenvalue errors in each iteration. Instead, we compare the Hausdorff distance between the final output $\varLambda_{h}^{\text{final}}$ (at the last iteration) with the discrete eigenvalue cluster $\varLambda_{h}^{(\ell)}$ (with two elements converging to the same number) found at the $\ell$th adaptive iteration. The curve exhibits an overall decreasing trend, even if not a monotonic decrease. The CL values at each iteration are compared to the CL value calculated from the final eigenvalue iterate in Figure 6(f), which illustrates convergence of the CL approximations.
5.2 A nested anti-resonant nodeless fiber
(a) Initial mesh for the NANF fiber.
(b) Final mesh for the NANF fiber after the adaptive algorithm.
@@ 5.2 A nested anti-resonant nodeless fiber
$10^{5.4}$$10^{5.6}$$10^{5.8}$$10^{6}$$10^{6.2}$$10^{-9}$$10^{-8}$$10^{-7}$$10^{-6}$$10^{-5}$umber of d.o.f.s iteration $\ell$$d(\varLambda_{h}^{\text{final}},\varLambda_{h}^{(\ell)})$
(e) History of adaptive convergence.
$10^{5.4}$$10^{5.6}$$10^{5.8}$$10^{6}$$10^{6.2}$$10^{-4}$$10^{-3}$umber of d.o.f.s at iteration $\ell$$\text{CL}_{\ell}$
(f) History of computed loss values.
Figure 7: Results from the adaptive algorithm applied to NANF.
@@
We employ the following parameters for the simulation. The length scale is $L=1.5\times 10^{-5}$ m; the operating wavelength is determined by $k=\frac{2\pi}{1.8}\times 10^{6}\leavevmode\nobreak\ \text{m}^{-1}$; the refractive index is $n_{\rm air}=1.00027717$ in the air regions, and $n_{\rm glass}=1.44087350$ in the glass ring. We utilize the following non-dimensionalized radii and thicknesses: the core of the fiber is a circle of radius $r_{\rm core}=1.0$; the outer capillary is determined by an outer radius $r_{\rm cap,\ o}=0.832$, and a thickness of $t_{\rm cap,\ o}=0.028$; the inner capillary is defined by an inner radius $r_{\rm cap,\ i}=0.4$, and a thickness of $t_{\rm cap,\ i}=0.028$; the cladding has an inner radius $r_{\rm clad}=1.33333333$, an outer radius $r_{\rm buffer}=2.0$, and a thickness $t_{\rm clad}=0.66666667$; the glass ring has an inner radius $r_{\rm inner}=2.692$, an outer radius $r_{\rm outer}=4.35866667$, and a thickness $t_{\rm clad}=1.66666667$; and the PML region starts at $r_{\rm outer}$ and ends on $r_{\rm PML}=7.02533333$. We consider a PML strength (see (14)) of value $\alpha=2.0$.
The results in Figure 7 confirm that fine-scale modal features exist for this fiber as well. The initial mesh (Figure 7(a)) is relatively coarse, with small elements used only to conform to the geometry of the thin glass capillaries and their melding into the cladding (as in the ARF case). We make the elements in the core region smaller (with $0.2\lesssim h\lesssim 0.5$), to capture the expected core-localized fundamental mode of this fiber and verify the saturation assumption. Again, the adaptive meshing algorithm found (in Figure 7(b)) that significant refinements were needed in the outer glass cladding; however, it also determined that some refinements were warranted in the hollow-core near larger capillaries that bound the core. The captured fine-scale modal features are visible in Figure 7(d). The convergence of the eigenvalues and CL values during the adaptive process are illustrated in Figures 7(e) and 7(f).
5.3 A photonic bandgap fiber
Finally, we consider the PBG microstructure design in Figure 1(d), previously studied in [24, 7, 19]. The lattice arrangement of the dielectric rods in the fiber can be modified to change the guiding properties of the fiber. We base our model parameters on the descriptions in [19].
@@ 5.3 A photonic bandgap fiber
(b) Final mesh for the PBG fiber after the adaptive algorithm.
@@
$10^{5}$$10^{6}$$10^{-11}$$10^{-8}$$10^{-5}$$10^{-2}$umber of d.o.f.s at iteration $\ell$$d(\varLambda_{h}^{\text{final}},\varLambda_{h}^{(\ell)})$
(e) History of adaptive convergence.
$10^{5}$$10^{6}$4,05040504{,}0504 , 0504,10041004{,}1004 , 1004,15041504{,}1504 , 150Number of d.o.f.s at iteration $\ell$$\text{CL}_{\ell}$
(f) History of computed loss values.
Figure 8: Results from the adaptive algorithm applied to the PBG fiber.
The scale of the PBG fiber is $L=2.88753\times 10^{-6}$ m. The wavelength we consider is $k=\frac{2\pi}{8.25}\times 10^{7}\text{m}^{-1}$. The refractive index of the cladding and the outer region is $n_{\rm cladding}=1.45$, and the refractive index of the tubes is $n_{\rm tube}=1.8$. We have used a lattice with six dielectric rods, each with a (dimensionless) radius of $r_{\rm tube}=0.57142857$, encircling a fiber core of radius $r_{\rm core}=1.0$. The inner radius of the PML region is $r_{\rm outer}=3.80952380$, and the outer radius of the PML region is $r_{\rm PML}=5.80952380$.
The results are in Figure 8. The initial mesh (Figure 8(a)) is relatively coarse, although we employed smaller elements in the core (with $h\approx 0.25$). For this fiber, as seen from Figure 8(b), the adaptive process targets refinement within the dielectric rods that surround the solid dielectric core. We see finer scale ripples within these rods being resolved in Figure 8(d). Again, these are better seen in the longitudinal component, and are hardly visible in the total intensity plot (in Figure 8(c)). The convergence of the eigenvalues during the adaptive process depicted in Figure 8(e) shows a monotonic descent, unlike what was found for the anti-resonant fibers (and like the prior cases, the discrete cluster $\varLambda_{h}^{(\ell)}$ here generally had two elements). The convergence of CL values is seen in Figure 8(f).